# Connect the Dots: Knowledge Graph–Guided Crawler Attack on Retrieval-Augmented Generation Systems

### v1

@@ 3.2 Adaptive Stochastic Coverage Problem
The Adaptive Stochastic Coverage Problem (ASCP) considers scenarios in which each action reveals information in a stochastic manner, and subsequent decisions depend on the observations made so far. In this setting, each action corresponds to selecting a set whose information coverage is uncertain until it is executed. The coverage outcome of the same item can vary depending on the results of previously chosen actions. The objective is to maximize the overall information coverage within limited budgets, such as a limited number of actions.
An ASCP sepcification can formally specified using item universe $\mathcal{U}$, action space $\mathcal{Q}$, stochastic query outcome $O(q)$, cost and budget $c(q),B$, and coverage function $f(S)$
  • Item universe ($\mathcal{U}$): The item universe $\mathcal{U}$ represents the set of all items that are available for coverage. These items are the targets for the actions, and the goal is to maximize coverage across this universe.
  • Action space ($\mathcal{Q}$): The action space $\mathcal{Q}$ consists of all possible actions that can be taken. Each action $q$ corresponds to covering a set of items, but the outcome of this action is uncertain until it is executed.
  • Stochastic query outcome ($O(q)$): The stochastic query outcome $O(q)$ defines the uncertain result of executing action $q$. It captures the probabilistic nature of the outcome, reflecting how the coverage outcome can vary depending on prior actions.
  • Cost and budget ($c(q),B$): The cost $c(q)$ represents the resources required to perform action $q$, and $B$ is the total available budget. The objective is to maximize coverage while staying within the budget constraint.
  • Coverage function ($f(S)$): The coverage function $f(S)$ quantifies the total information coverage achieved by a set of actions $S$. It reflects the degree to which the selected actions have covered items from the item universe $\mathcal{U}$.
@@
We reduce the data extraction attack against RAG systems, which we term RAG Crawling to the ASCP to guide our solution (see Sec. 5).
@@ 4 Threat Model
Scenario. As illustrated in Fig. 3, we consider a production RAG service accessed through a black box interface such as an API or chat. Internally, a retriever selects passages or documents from a private corpus $\mathcal{D}$ and a generator composes an answer from the query and the retrieved evidence [karpukhin2020dense, lewis2020retrieval]. In deployment the generator may rewrite, summarize, or refuse according to guardrails and access policies [multiquery, beck2025ragqr]. The service does not reveal retrieval scores or document identifiers. The attacker has no insider access and interacts as a normal user. The attacker aims to extract as much of $\mathcal{D}$ as possible while remaining unnoticed by steering retrieval to cover new items across turns.
Prior work has identified two broad classes of such extraction attacks in RAG systems: Explicit attacks append overt adversarial instructions to elicit verbatim disclosure of retrieved text [cohen2024unleashing, jiang2024rag, di2024pirates, zeng2024good]. Implicit attacks use benign-looking queries and aggregate factual content that the system legitimately returns across turns, reconstructing private material via paraphrases and fragments [wang2025silent]. Our focus is the implicit regime, which better matches real usage and evades straightforward refusal rules.
Figure 3: Threat model of RAG Crawling. An external attacker interacts with a production RAG system only via the public interface. For each query, the retriever selects a set of documents from the private corpus and the LLM generates an answer conditioned on this evidence. Across turns, the attacker aims to stealthily expand the union of leaked retrieved items under a query budget.
Adversary’s Goal. The attacker aims to exfiltrate the private knowledge base $\mathcal{D}$ as completely as possible while avoiding detection. Leakage may appear as verbatim text, paraphrases, summaries, or factual statements that reflect retrieved evidence. For each round $t$, the adversary submits a query $q_{t}$. In response, the RAG system internally consults a hidden set of documents $\mathcal{D}_{k}(q_{t})\subseteq\mathcal{D}$ to generate its answer. The adversary’s goal is to maximize the number of unique documents retrieved over a sequence of $B$ queries, thereby learning about the system’s underlying knowledge base. This budget $B$ models real-world constraints such as API costs and the need to evade detection. This objective is formally expressed as the following:
@@
Stealth remains a parallel objective: each $q_{t}$ should look natural and innocuous to avoid detection or refusal.
Adversary’s Capabilities. We assume a black box setting. The attacker submits adaptive natural language queries $q_{t}$ and observes only final answers $a_{t}$. They cannot see retrieval scores, document identifiers, or intermediate states, and they cannot read internal indices or modify retriever weights. They do not poison the target corpus and do not use privileged operator tools. Following prior work [cohen2024unleashing, jiang2024rag, wang2025silent], the attacker knows only the topic of the target knowledge base, such as a keyword or short phrase that can be inferred from public materials or light probing. Under these constraints, leverage comes solely from crafting query sequences and interpreting observed answers.
@@ 5 Reducing RAG Crawling to ASCP
We now explain how we model the data extraction in RAG, RAG Crawling, as ASCP (introduced in Sec. 3.2). This formal mapping clarifies how we can potentially apply the greedy coverage-maximizing approach in our setting:
  • Item universe ($\mathcal{U}$): We take $\mathcal{U}$ to be the set of atomic documents or passages in the private corpus $\mathcal{D}$. Essentially, the “items” to be covered are the documents in $\mathcal{D}$.
  • Action space ($\mathcal{Q}$): We equate the set of possible actions $\mathcal{Q}$ with the set of all valid queries the attacker could ask might pose. In theory $\mathcal{Q}$ is infinite (any string could be a query), but we will address how to search it later. For now, conceptually, each $q\in\mathcal{Q}$ is a possible question the attacker can pose to the target system.
  • Stochastic query outcome ($\Phi$): When the attacker issues a particular query $q$, the RAG system will retrieve some set of documents $\mathcal{D}_{k}(q)\subset\mathcal{D}$ (Sec. 3.1) and then produce an answer. We model this retrieval result as a random subset drawn from a distribution $\Phi(\cdot\mid q)$. Essentially, $\Phi$ encapsulates our uncertainty about what part of $\mathcal{D}$ will be exposed by query $q$.
  • Cost and budget: We assume each query has a unit cost (e.g., each question counts as 1 against the budget $B$). This is a reasonable simplification: perhaps some queries could be considered “more costly” if they are longer or trigger more computation, but in our threat model the primary cost is just the opportunity cost of using up one of the limited queries we can send. Thus, a budget $B$ simply means the attacker is limited to $B$ queries in total.
  • Coverage function: We define $f(S,\Phi)$ as the fraction of unique documents in $\mathcal{D}$ that have been retrieved at least once by the set of queries $S$. Since the attacker cannot directly observe $\mathcal{D}_{k}(q)$, this is a latent coverage measure.
Under this reduction, the attacker is effectively an adaptive agent who, at each step $t=1,2,\dots,B$, selects a query $q_{t}\in\mathcal{Q}$ based on the interaction history so far, aiming to maximize the overall expected coverage $f(S,\Phi)$ by the end of $B$ queries.
We now define the notion of adaptive marginal gain under the RAG Crawling model and show that it satisfies the core assumptions required for approximate greedy optimization in ASCP.
Definition 1 (Conditional Marginal Gain in RAG Crawling).
Let $S=\{q_{1},\dots,q_{t}\}$ denote the set of queries issued so far, and let $\psi$ denote the observable interaction history, that is, the sequence of issued queries and their corresponding observable system responses such as generated answers. Let $\Phi$ represent the underlying stochastic retrieval process that maps queries to latent sets of retrieved documents. The conditional marginal gain of issuing a new query $q\in\mathcal{Q}$ given history $\psi$ is defined as
$\Delta(q\mid\psi):=\mathbb{E}_{\Phi}\left[f(S\cup\{q\},\Phi)-f(S,\Phi)\;\middle|\;\psi\right],$ (4)
@@
The proof can be found in Appendix 9.2. Building on Theorem 1 and Theorem 2, we have the following:
Theorem 3 (Near-optimality guarantee of RAGCrawler).
@@
$q_{t}\in\arg\max_{q\in\mathcal{Q}}\Delta(q\mid\psi_{t-1}),$ (5)
achieves an approximately optimal expected coverage.
While this theorem provides a principled optimization target, implementing an adaptive greedy strategy in practice is far from trivial. Several challenges arise when translating the theoretical ASCP framework into a practical attack against real-world RAG systems:
  • The CMG is unobservable. The attacker cannot directly observe the true coverage gain of a query, as the retrieved document sets remain hidden and inaccessible.
  • The attacker’s action space is intractable. The action space $\mathcal{Q}$ includes all valid natural language strings, making it effectively infinite. Without additional structure or priors, exhaustive search for high-gain queries becomes computationally infeasible.
  • Real-world RAG systems’s defense. Queries must appear natural and comply with safety policies to avoid refusal, rejection, or detection by safety filters.
In next section, we explain how RAGCrawler, a black-box crawling framework designed, reconstructs the latent knowledge base of a RAG system under a fixed query budget. Grounded in the Adaptive Stochastic Coverage Problem (ASCP) formulation, RAGCrawler approximates the theoretical greedy coverage-maximization strategy by maintaining an evolving knowledge-graph state, identifying high-gain anchors in the semantic space of this state, and realizing them through stealthy queries.
@@ 5.1 Overview
In this section, we explain the operational workflow of RAGCrawler, illustrated in Fig. 4. The attacker begins by issuing an initial query $q_{0}$ to the victim RAG system and receiving the corresponding response $a_{0}$. Upon receiving a query, the RAG retrieves a hidden set of documents from its private corpus and synthesizes an answer based on the retrieved evidence. Although the retrieved content is not directly observable, the generated response inevitably reveals partial fragments of the underlying corpus. The KG-Constructor module parses $a_{0}$ into structured triples and integrates them into the initial attacker-side KG $\mathcal{G}_{0}$, which serves as an observable approximation of the latent corpus coverage. Based on this state, the Strategy Scheduler estimates the CMG over candidate entity–relation anchors in $\mathcal{G}_{0}$, selects the most promising anchor $Anc_{0}$, and forwards it to the Query Generator. The Query Generator realizes this anchor as a fluent and policy-compliant query $q_{1}$, which is then sent to the RAG system to obtain the next response $a_{1}$. The new answer $a_{1}$ is subsequently processed by the KG-Constructor to update the knowledge graph to $\mathcal{G}_{1}$. Generalizing this process, at iteration $t$ the attacker issues $q_{t}$, receives $a_{t}$, and updates the state $\mathcal{G}_{t}$ through the same sequence of modules, forming a closed adaptive loop of state estimation, strategic planning, and query realization.
@@
Figure 4: The workflow of RAGCrawler. At each step, (1) The KG-Constructor processes the latest system response to update the knowledge graph. (2) The Strategy Scheduler analyzes this graph to select a strategic anchor. (3) The Query Generator then uses this anchor to formulate the next query, completing the loop.
@@ 5.2 KG-Constructor
This module addresses the first challenge in the ASCP formulation: the CMG of the attacker’s action is unobservable. Since the attacker cannot directly observe which documents have been retrieved by the RAG system, the true coverage increment of each query remains hidden. The goal of this module is therefore to construct and maintain an evolving KG $\mathcal{G}_{t}$ that approximates the latent coverage function $f(S^{\pi},\Phi)$, making CMG estimable from surface-level answers. The workflow is shown in Fig. 5.
@@
Why existing approaches do not work. Existing KG construction methods are ill-suited for our dynamic setting. Traditional techniques such as OpenIE and schema-constrained extraction [yates2007textrunner, stanovsky2018supervised], rely on fixed relation inventories and curated ontologies, assuming stable data access and consistent syntax. Recent LLM-based approaches [papaluca2024zero, bian2025llm] show promise for adaptive graph construction but conflicts with the need for efficient incremental updates that do not require full access to the underlying data. Furthermore, graph-augmented RAG frameworks such as GraphRAG [edge2024local] and LightRAG [guo2024lightrag] use LLMs to extract triples from documents within the corpus and leverage the graph to enhance retrieval. These frameworks primarily focus on optimizing retrieval quality by merging exactly identical entities and relations. While this improves retrieval efficiency, it neglects the need for compact and efficient coverage tracking across multiple rounds of dynamic query responses.
Our methods. We develop a KG-Constructor that emphasizes efficient incremental updates. Instead of requiring access to the entire dataset or repeatedly rebuilding the global graph, we maintain the KG in a lightweight manner by constructing a subgraph at each step and merging it into the existing graph. As illustrated in Fig. 5, the workflow consists of three main modules:
1) Topic-Specific Prior. We begin by establishing a topic-specific prior that defines the contextual focus of the constructed KG. This prior acts as a soft constraint on the LLM’s extraction process, guiding it toward topic-relevant entities and relations while filtering out off-topic or noisy content. By narrowing the semantic search space, it enables more targeted and consistent knowledge acquisition. Concretely, we prompt the LLM to infer a compact set of entity categories and relation types that characterize the target domain. For example, in a medical corpus, entity categories may include disease, symptom, and treatment, while relation types may include has_symptom, treated_by, and caused_by. During extraction, the LLM is instructed to prioritize these schema elements when forming structured triples, ensuring that all generated relations remain within the topical boundary. This bounded, schema-guided design offers two key advantages. First, it stabilizes relation typing and prevents the introduction of spurious or semantically inconsistent edges, maintaining structural coherence across incremental updates. Second, it keeps the prompt context compact and computationally efficient: the LLM does not need to reference all existing nodes and relations in the graph, but only the abstract schema of entity and relation types as contextual constraints. This design greatly reduces prompting overhead, enabling scalable LLM calls while preserving high topical precision in the extracted knowledge.
@@
Figure 5: The workflow of KG-Constructor. Guided by a Topic-Specific Prior, an Iterative Extraction and Reflection process generates a knowledge subgraph ($\Delta\mathcal{G}_{t}$), which is then refined through Incremental Graph Update and Semantic Merging to produce the final graph state, $\mathcal{G}_{t}$.
@@
3) Incremental Graph Update and Semantic Merging. After obtaining the new knowledge subgraph $\Delta\mathcal{G}_{t}$, it is incrementally integrated with the existing graph $\mathcal{G}_{t-1}$ to form the updated graph $\mathcal{G}_{t}$. This integration is implemented as a lightweight merging operation without rebuilding graph from scratch. To ensure that the graph structure reflects genuine semantic expansion rather than redundant surface-level variations, we perform a semantic merging stage offline. All entity mentions and relation phrases are encoded into a shared embedding space using a pre-trained encoder. Node or edge pairs whose cosine similarity exceeds a threshold $\tau$ are automatically identified and merged. This joint reconciliation of entities and relations consolidates semantically equivalent concepts, normalizes synonymy across updates, and maintains a compact yet expressive knowledge graph. This provides a stable basis for CMG estimation and planing space for Strategy Scheduler (Sec. 5.3).

### v2

@@ 3.2 Adaptive Stochastic Coverage Problem
The Adaptive Stochastic Coverage Problem (ASCP) models budgeted coverage maximization in sequential decision-making under uncertainty. At each step, an agent selects an action, observes what it reveals, and adapts future actions based on the observations so far. The objective is to maximize expected coverage under a limited budget, such as a limited number of actions. An ASCP instance can be specified by:
  • Item universe ($\mathcal{U}$): The item universe $\mathcal{U}$ represents the set of all items that are available for coverage.
  • Action space ($\mathcal{Q}$): The action space $\mathcal{Q}$ consists of all possible actions. Each action $q$ can potentially reveal (cover) a subset of items in $\mathcal{U}$.
  • Stochastic outcome (realization $\Phi$ and $O_{\Phi}(q)$): Uncertainty is modeled by a latent realization $\Phi$ (a “state of the world”) drawn from a distribution. Conditioned on $\Phi$, executing action $q$ reveals an outcome set $O_{\Phi}(q)\subseteq\mathcal{U}$.
  • Cost and budget ($c(q),B$): The cost $c(q)$ represents the resources required to perform action $q$, and $B$ is the total available budget.
  • Coverage function ($f(S,\Phi)$): The coverage function $f(S,\Phi)$ quantifies the utility of what has been revealed by a set of actions $S$ under realization $\Phi$. For coverage objectives, $f$ is commonly a function of the union $\bigcup_{q\in S}O_{\Phi}(q)$.
@@
We reduce RAG knowledge-base stealing to ASCP by formalizing it as RAG Crawling (Sec. 5), enabling CMG-based greedy optimization as a principled strategy for globally planned stealing under a query budget.
@@ 4 Threat Model
Scenario. As illustrated in Fig. 3, we consider a production RAG system accessed through a black box interface such as an API or chat. Internally, a retriever selects passages or documents from a corpus $\mathcal{D}$ and a generator composes an answer from the query and the retrieved documents [karpukhin2020dense, lewis2020retrieval]. In deployment the generator may rewrite, summarize, or refuse according to guardrails and access policies [multiquery, beck2025ragqr]. The service does not reveal retrieval scores or document identifiers. The attacker has no insider access and interacts as a normal user. The attacker aims to steal as much of $\mathcal{D}$ as possible by steering retrieval to cover new items across turns.
Prior work has identified two classes of such stealing attacks in RAG systems: Explicit attacks append overt adversarial instructions to elicit verbatim disclosure of retrieved text [cohen2024unleashing, jiang2024rag, di2024pirates, zeng2024good]. Implicit attacks use benign-looking queries and aggregate factual content that the system legitimately returns across turns, reconstructing corpus via paraphrases and fragments [wang2025silent]. Our focus is the implicit regime, which better matches real usage and evades refusal rules.
Figure 3: Threat model. An external attacker has black-box access to a deployed RAG system and can only interact via its public query interface. For each query, the retriever selects a set of documents from the corpus and the LLM generates an answer conditioned on these documents. Across turns, the attacker aims to stealthily expand the union of leaked retrieved items under a query budget.
Adversary’s Goal. The attacker aims to steal the knowledge base $\mathcal{D}$ as completely as possible while remaining stealthy. Leakage may appear as verbatim text, paraphrases, summaries, or factual statements grounded in retrieved documents. At round $t$, the adversary submits $q_{t}$ and the system internally consults a hidden retrieved set $\mathcal{D}_{k}(q_{t})\subseteq\mathcal{D}$ to generate the answer. Over $B$ queries, the objective is to maximize the fraction of unique retrieved items:
@@
The budget $B$ models practical constraints such as API cost and the need to limit suspicious activity.
Adversary’s Capabilities. We assume black box access: the attacker submits adaptive natural language queries and observes only final answers. They have no access to retrieval scores, identifiers, indices, or intermediate states, and they cannot modify the retriever or poison the corpus. Following prior work [cohen2024unleashing, jiang2024rag, wang2025silent], we assume a coarse topic phrase of the target knowledge base is known, as it is often exposed by product descriptions, onboarding pages, or brief interaction. We also evaluate a weaker attacker with no topic prior that issues one benign probe query and compresses the response into a short topic phrase using an attacker-side LLM (the probe counts toward the budget; Appendix D).
@@ 5 Reducing RAG Crawling to ASCP
We formalize the attacker’s goal as RAG Crawling and reduce it to ASCP (Sec. 3.2). In this way, coverage-maximizing adaptive greedy policies can provide a principled benchmark. We instantiate the ASCP components as follows:
  • Item universe ($\mathcal{U}$). We set $\mathcal{U}:=\mathcal{D}$, the attacker-hidden corpus, where each item is an atomic document or passage.
  • Action space ($\mathcal{Q}$). We set $\mathcal{Q}$ to be the set of all valid natural-language queries the attacker may issue to the victim RAG.
  • Stochastic outcome (realization $\Phi$ and $O_{\Phi}(q)$). Let $\Phi$ denote a latent realization of the victim RAG pipeline that determines retrieval outcomes (including any nondeterminism and attacker-unknown processing such as rewriting or multi-query retrieval). Conditioned on $\Phi$, issuing query $q$ leads the pipeline to use a hidden set of $k$ retrieved items to generate the final answer; we denote this outcome by
    $O_{\Phi}(q):=\mathcal{D}^{\Phi}_{k}(q)\subseteq\mathcal{D}.$ (4)
  • Cost and budget ($c(q),B$). We assume unit query cost $c(q)=1$, so the attacker can issue at most $B$ queries. This matches common deployment settings where the service API charges on a per-query basis: each query consumes one billable request regardless of its exact content.
  • Coverage function ($f(S,\Phi)$). For a set of queries $S\subseteq\mathcal{Q}$ and realization $\Phi$, the coverage function is
    $f(S,\Phi):=\left|\bigcup_{q\in S}O_{\Phi}(q)\right|\,$, (5)
    which matches adversary’s goal (Eq. 3) up to a constant.
We next define conditional expected marginal gain (CMG) as follows. CMG is the oracle quantity optimized by adaptive greedy in ASCP.
Definition 1 (Conditional Expected Marginal Gain in RAG Crawling).
Let $\psi^{R}_{t-1}$ denote the current retrieval-level partial realization (i.e., the executed query–outcome pairs under the latent realization $\Phi$) after t1t\!-\!1 issued queries, and let $S_{t-1}\subseteq\mathcal{Q}$ be the set of queries issued so far. The conditional expected marginal gain (CMG) of a new query $q\in\mathcal{Q}$ given $\psi^{R}_{t-1}$ is
$\Delta(q\mid\psi^{R}_{t-1}):=\mathbb{E}_{\Phi\mid\psi^{R}_{t-1}}\!\left[f(S_{t-1}\cup\{q\},\Phi)-f(S_{t-1},\Phi)\right].$ (6)
We then show that RAG Crawling satisfies the assumptions required by ASCP’s greedy approximation theory.
@@
The proof can be found in Appendix A. Building on Theorem 1 and Theorem 2, we have the following:
Theorem 3 (Near-optimality guarantee of the adaptive greedy policy for RAG Crawling).
@@
$q_{t}\in\arg\max_{q\in\mathcal{Q}}\Delta(q\mid\psi^{R}_{t-1}),$ (7)
achieves a (11/e)(1-1/e)-approximation to the optimal expected coverage under the same query budget.
Figure 4: The workflow of RAGCrawler. At each step, (1) The KG-Constructor processes the latest system response to update the knowledge graph. (2) The Strategy Scheduler analyzes this graph to select a strategic anchor. (3) The Query Generator uses this anchor to formulate the next query.
6 RAGCrawler
Theorem 3 provides a theoretical guarantee on optimal coverage of RAG Crawling. We now present RAGCrawler, which instantiates the adaptive greedy policy for RAG knowledge-base stealing and enables real-world attacks that achieve 66.8% coverage on average with only 1,000 queries.
@@ 6.1 Overview
In this section, we explain the workflow of RAGCrawler (Fig. 4). RAGCrawler implements a closed-loop crawler that alternates between graph state estimation, anchor scheduling, and natural query realization.
At round $t$, the attacker issues $q_{t}$ and observes only the answer text $a_{t}$, while the retrieved documents remain hidden in the black-box setting. Each answer $a_{t}$ is processed by KG-Constructor to extract a step-wise knowledge subgraph and update the attacker-side state to $\mathcal{G}_{t}$. Based on $\mathcal{G}_{t}$, the Strategy Scheduler selects a semantic anchor $Anc_{t}$ to approximate which direction is likely to yield high marginal coverage gain in subsequent queries, where we define $Anc_{t}\triangleq(e_{t}^{*},r_{t}^{*})$ as an entity with an optional relation target. Finally, the Query Generator realizes $Anc_{t}$ into a fluent and policy-compliant query $q_{t+1}$, sends it to the victim RAG, and the loop repeats. To bootstrap a non-empty graph state, we run a brief topic conditioned seeding phase that generates a few broad queries, and all seeds count toward the budget.
@@ 6.2 KG-Constructor
This module addresses the first challenge in the ASCP formulation: the CMG of the attacker’s action is unobservable. Since the attacker cannot directly observe which documents have been retrieved by the RAG system, the true coverage increment of each query remains hidden.
KG-Constructor makes coverage gains estimable by converting answer text into an attacker-side knowledge graph that compactly summarizes revealed content. Specifically, we maintain an evolving graph state $\mathcal{G}_{t}$ that approximates the latent coverage function $f(S,\Phi)$, making CMG estimable from surface-level answers. The workflow is shown in Fig. 5.
@@
Why existing approaches do not work. Existing KG construction pipelines do not fit our setting because we require consistent, lightweight incremental updates from partial and evolving answers, without access to the underlying corpus. OpenIE-style extraction [yates2007textrunner] produces free-form relations that drift across rounds, causing redundancy and unstable typing, while schema-constrained methods [stanovsky2018supervised] reduce variance but typically assume a fixed ontology and full-data access. LLM-based construction [papaluca2024zero, bian2025llm] and graph-augmented RAG frameworks (e.g., GraphRAG [edge2024local], LightRAG [guo2024lightrag]) mainly target one-shot document parsing and retrieval quality, and thus do not provide compact, efficient coverage tracking over many rounds of dynamic responses.
@@
Figure 5: KG-Constructor. Guided by a Topic-Specific Prior, the iterative extraction and reflection process generates a knowledge subgraph ($\Delta\mathcal{G}_{t}$), which is then refined through incremental graph update to produce the final graph state $\mathcal{G}_{t}$.
Our methods. Our KG-Constructor supports efficient incremental maintenance by extracting a per-answer subgraph and merging it into the global graph. At round $t$, the attacker treats the observed answer $a_{t}$ as the only evidence and extracts a typed triple set $\{(h,r,t)\}$, where $h$ and $t$ are entity mentions (or attribute values) and $r$ is a semantic relation type. We build a knowledge delta $\Delta\mathcal{G}_{t}$ using unique $h/t$ mentions as nodes and triples as directed edges, and merge it into $\mathcal{G}_{t-1}$ to obtain $\mathcal{G}_{t}$ without rebuilding from scratch. As shown in Fig. 5, the workflow has three modules:
1) Topic-Specific Prior. We derive a lightweight topic-specific schema prior to keep extraction focused and consistent across rounds. Given only a coarse topic signal, we instruct an attacker-side LLM to propose a compact set of entity categories and relation types for the domain (e.g., disease, symptom, treatment, and relations such as has_symptom, treated_by). During extraction, the LLM prioritizes these schema elements as soft constraints, filtering off-topic content and keeping triples within the topical boundary. This schema stabilizes relation typing across rounds and keeps prompts compact because the LLM conditions on the abstract schema rather than the full graph, reducing overhead while preserving topical precision.
@@
3) Incremental Graph Update and Semantic Merging. We finally update the global graph by merging $\Delta\mathcal{G}_{t}$ into $\mathcal{G}_{t-1}$ with a lightweight attacker-side operation. This integration is implemented as a lightweight merging operation without rebuilding graph from scratch. To ensure that the graph structure reflects genuine semantic expansion rather than redundant surface-level variations, we perform a semantic merging stage. All entity mentions and relation phrases are encoded into a shared embedding space using a pre-trained encoder. Node or edge pairs whose cosine similarity exceeds a threshold $\tau$ are automatically identified and merged.
This normalization consolidates synonymous mentions across rounds, keeps $\mathcal{G}_{t}$ compact and stable for CMG estimation and planning space for Strategy Scheduler (Sec. 6.3).